OpenStudy (anonymous):

Milk is leaking from a carton at a rate of 3ml/min. There is 1.2L of milk in the carton at 11:00am. How much will be left in the carton at 1:00pm

OpenStudy (anonymous):

Oh I get it now thank you!

OpenStudy (anonymous):

\[1.2L−2hours×\frac{ 3ml }{ 1\min }×\frac{ 60\min }{ 1hour }×\frac{ 1L }{ 1000ml }\]